Q A couple of days after the Christmas holiday, you received a visit from the City Health Department. The official was interested in learning if you had attended your neighborhood Christmas party. You asserted “Yes” and explained that you enjoyed the celebratory festivity. The official proceeded with asking questions if you had experienced vomiting, fever, rash on your skin and headache 24 to 48 hours after the party. Additionally, he asked if you had consumed food during the block party. Finally, he asked about the duration of your stay at the party. He specifically wanted to know if you had stay 2 hours or more than 2 hours. You answered all the questions as best as you could recall. Finally, at the closing of the interview, you stated that in future you would be a healthcare provider and you were interested in assisting with his analyses and interpretation of the findings. About a week later you received the following information. (40 points)

Table 1: Distribution of new year party goers by type of food consumed

Turkey

Eggnog

Turkey

126

221

Eggnog

221

99

Table 2: Distribution of party goers who became ill

Rash

Vomit

Fever

Diarrhea

10

50

40

Vomit

50

80

30

Fever

40

30

0

A: How many attended the picnic?

B: What is the proportion of party goers who consumed only Turkey, only Eggnog, Turkey and Eggnog?

C: What is the proportion of party goers who experienced rash only, vomiting only, Fever only,

D: What is the proportion of party goes who experienced at least two symptoms?

E: What is the ratio of individuals who experienced rash only to those who experienced vomiting only?
Answer #1

A: How many attended the picnic?

126 + 21 + 99 = 446

B: What is the proportion of party goers who consumed only Turkey, only Eggnog, Turkey and Eggnog?

P(only Turkey) = 126 126 446 126 + 221499

P(only Eggnog) = 199 126 + 221499 199 446

P(Turkey and Eggnog) = 221 126 + 221499 221 446

C: What is the proportion of party goers who experienced rash only, vomiting only, Fever only,

P(rash only) = 10+ 50+40+ 50 + 80 +30+40+30 +0 330

P(vomiting only) = 80 10+ 50+40+ 50 + 80 +30+40+30 +0 | 80 330

P(Fever only) = 10+ 50+40+ 50 + 80 +30+40+30 +0 330

D: What is the proportion of party goes who experienced at least two symptoms?

-1 = _0+ 0 + 0 + 0 + 0 + 01 + 08 + 08 + 0 0 +08 +0 08

E: What is the ratio of individuals who experienced rash only to those who experienced vomiting only?

\frac{0}{80}

Please give me a thumbs-up if this helps you out. Thank you!